OpenStudy (anonymous):

Let f(x) = x2 − 3x − 7. Find f(−3).

OpenStudy (texaschic101):

f(x) = x^2 - 3x - 7 sub in -3 for x f(-3) = -3^2 - 3(-3) - 7 can you do the math and finish ??

OpenStudy (anonymous):

i just solve right

OpenStudy (skullpatrol):

It is actually called "evaluate" because you are asked to find a "value."

OpenStudy (skullpatrol):